# Nernst effect and its thickness dependence in superconducting NbN films

### v2

@@ II Samples and methods
The properties of the five samples studied in this work are summarized in table 1, and the temperature dependence of their electrical resistivity is shown in Figure 1. In all samples, resistivity increases with decreasing temperature, as expected in a dirty metal. As the thickness is reduced, resistivity increases from 70 $\rm\mu\Omega.cm$ at 30 nm to 580 $\rm\mu\Omega.cm$ at 4 nm. Note that while the resistivity increases by a factor 8, the sheet resistance increases by a factor $>$50. Concomitantly, the critical temperature decreases from 15.5 K to 8.9 K. Both these features were reported in previous studies on NbN thin films. They are generic to disordered superconducting thin films [47, 48, 49]. Even in our thinnest sample is not an insulator. Its sheet resistance is well below $h/4e^{2}=6.45\rm k\Omega$, and the normal state resistivity displays a mild temperature dependence, characteristic of weak localization in a dirty metal.
@@ III Results
Table 1: Properties of the five NbN thin films used in this study. $\rho_{max}$ is defined as the maximum of resistivity above the superconducting transition. The residual resistivity ratio RRR is defined as $\rho$(300K)/$\rho_{max}$
@@ III.1 The upper critical field
Fig 2f displays how the critical temperature changes as a function of the magnetic field. We used Werthamer–Helfand–Hohenberg theory [50] and our $H_{c2}(T)$ data to estimate the zero-temperature upper critical field, $H_{c2}(0)$.
@@
The results are shown table 1. Values of coherence obtained with this method will be wwritten $\xi_{0}^{Hc2}$. Remarkably, despite an almost twofold variation of the critical temperature, the zero-temperature coherence length is almost the same. Note that for all studied samples, $\xi_{0}\sim$ 4.5 $\pm$ 0.3 nm is larger or close to the sample thickness.
@@ III.2 Concomitance between the Nernst response and the electric resistivity
Above the critical temperature, the amplitude of $N$ decreases drastically, but remains finite. Note also that the peak Nernst effect decreases with increasing thickness, and therefore with the amplitude of the resistivity. It decreases from 2rm $\rm\mu V/K$ in the 12.5 nm sample to 1.3 $\rm\mu V/K$ in the 20 nm one. This is a $\approx$ 1.54 times decrease, comparable to the $\approx$ 1.6 decrease in resistivity. We will discuss this in more detail below. Note, however, that the order of magnitude of this vortex Nernst peak is a few $\mu V/K$ in all samples, and comparable to what was found in many superconductors with critical temperature varying from a few hundreds of mK to 100 K. This is in sharp contrast with metals, where the quasi-particle Nernst response can have an amplitude ranging from nV/K to a few mV/K [1].
@@ III.4 The Nernst coefficient in the normal state
Comparing our Fig 5, with Figure 8 in ref. [1], one can see that in our data the low-temperature regime (in which $\nu$ becomes field independent) is clearly visible. On the other hand, the high-field regime (where $\nu$ becomes temperature-independent) is not attained. It requires a magnetic field significantly larger than 10 T. Nevertheless, a tendency towards it is detectable.
@@ III.5 The amplitude and the temperature dependence of the Nernst conductivity
The combination of resistivity and Nernst data allows us to determine the amplitude of $\alpha_{xy}$ in the samples. Fig 6 shows $\alpha_{xy}$ (that is, $\frac{N}{\rho}$) in three (4, 12.5 and 20 nm) samples at 0.5 T, as a function of the reduced temperature, $\varepsilon$=(T-Tc)/Tc. Here, $T_{c}$ is taken as the midpoint of the resistive superconducting transition. A field of 0.5 T is enough to reach the low field regime discussed previously, and $\nu$ only depends on the temperature. Remarkably, in a significant range of reduced temperature ($10.05<\varepsilon<$ 1), the three curves fall on one another. Using the 2D definition of $\alpha^{2D}_{xy}=N/R_{S}$, expressed in A/K, would have yielded a sample dependent Nernst conductivity, varying from 0.04 A/K at 4 nm to 0.2 A/K at 20 nm, a five fold increase. As seen in the figure, in this temperature range, which extends to twice the critical temperature, $\alpha_{xy}\propto\varepsilon^{-1}$. Given that $\xi^{2}\propto\varepsilon^{-1}$, this is, indeed, what is expected by Equation 4.
@@ IV.1 Dimensionality of the Gaussian fluctuations
In both equations, the Nernst conductivity is proportional to the quantum of thermoelectric conductance, $k_{b}e/\hbar$, and the only sample dependant parameter is the coherence length. In both equations, the temperature dependence is set by the ratio of $\xi$ and $\ell_{B}$. However, the difference in exponents (reflecting the difference in units (A/K in 2D, and A/K.m in 3D), leads to a qualitatively different temperature dependence: $\alpha_{xy}^{2D}\propto\varepsilon^{-1}$ and $\alpha_{xy}^{3D}\propto\varepsilon^{-1/2}$.
@@
Faced with this apparent paradox, we note first that our situation is neither two-dimensional (the thickness exceeds the coherence length), nor truly three-dimensional (there is a confinement along the c-axis). Now, as seen in equation 3, $\alpha_{xy}$ is the product of the Nernst signal, (which does not depend on dimensionality), and conductivity (which does). The theoretical 3D expression (Equation 7) is derived assuming a stack of 2D layers with $2\xi$ thick. It is the temperature dependence of $2\xi$, which changes the exponent of the power law. Assuming that the relevant length scale along the third direction is $2\xi_{0}$, which has no temperature dependence, Equation 4 becomes :

### v3

@@ II Samples and methods
The properties of the five samples studied in this work are summarized in table 1, and the temperature dependence of their electrical resistivity is shown in Figure 1. In all samples, resistivity increases with decreasing temperature, as expected in a dirty metal. As the thickness is reduced, resistivity increases from 70 $\rm\mu\Omega.cm$ at 30 nm to 580 $\rm\mu\Omega.cm$ at 4 nm. Note that while the resistivity increases by a factor 8, the sheet resistance increases by a factor >>50. Concomitantly, the critical temperature decreases from 15.5 K to 8.9 K. Both these features were reported in previous studies on NbN thin films. They are generic to disordered superconducting thin films [47, 48, 49]. Even in our thinnest sample is not an insulator. Its sheet resistance is well below $h/4e^{2}=6.45\rm k\Omega$, and the normal state resistivity displays a mild temperature dependence, characteristic of weak localization in a dirty metal.
@@ III Results
[…]
@@ III.1 The upper critical field
Fig 2f displays how the critical temperature changes as a function of the magnetic field. We used Werthamer–Helfand–Hohenberg theory [50] and our $H_{c2}(T)$ data to estimate the zero-temperature upper critical field, $H_{c2}(0)$ :
@@
The results are shown table 1. Values of coherence obtained with this method will be written $\xi_{0}^{Hc2}$. Remarkably, despite an almost twofold variation of the critical temperature, the zero-temperature coherence length is almost the same. Note that for all studied samples, $\xi_{0}\sim$ 4.5 $\pm$ 0.3 nm, that is shorter than the thickness of all samples, save for the thinnest one.
@@ III.2 Concomitance between the Nernst response and the electric resistivity
Above the critical temperature, the amplitude of $N$ decreases drastically, but remains finite. Note also that the peak Nernst effect decreases with increasing thickness, and therefore with the amplitude of the resistivity. It decreases from 2 $\rm\mu V/K$ in the 12.5 nm sample to 1.3 $\rm\mu V/K$ in the 20 nm one. This is a $\approx$ 1.54 times decrease, comparable to the $\approx$ 1.6 decrease in resistivity. We will discuss this in more detail below. Note, however, that the order of magnitude of this vortex Nernst peak is a few $\mu V/K$ in all samples, and comparable to what was found in many superconductors with critical temperature varying from a few hundreds of mK to 100 K. This is in sharp contrast with metals, where the quasi-particle Nernst response can have an amplitude ranging from nV/K to a few mV/K [1].
@@ III.4 The Nernst coefficient in the normal state
Comparing our Fig 5, with Figure 8 in ref. [1], one can detect in our data a low-field regime, in which $\nu$ becomes independent of magnetic field. On the other hand, the high-field regime (where $\nu$ becomes temperature-independent) is not attained in our measurements. It requires a magnetic field significantly larger than 10 T. Nevertheless, a tendency towards it is detectable.
@@ III.5 The amplitude and the temperature dependence of the Nernst conductivity
The combination of resistivity and Nernst data allows us to determine the amplitude of $\alpha_{xy}$ in the samples. Fig 6 shows $\alpha_{xy}$ (that is, $\frac{N}{\rho}$) in three (4, 12.5 and 20 nm) samples at 0.5 T, as a function of the reduced temperature, $\varepsilon$=(T-Tc)/Tc. Here, $T_{c}$ is taken as the midpoint of the resistive superconducting transition. A field of 0.5 T is enough to reach the low field regime discussed previously, and $\nu$ only depends on the temperature. Remarkably, in a significant range of reduced temperature ($0.05<\varepsilon<1$), the three curves fall on one another. Using the 2D definition of $\alpha^{2D}_{xy}=N/R_{S}$, expressed in A/K, would have yielded a sample dependent Nernst conductivity, varying from 0.04 A/K at 4 nm to 0.2 A/K at 20 nm, a five fold increase. As seen in the figure, in this temperature range, which extends to twice the critical temperature, $\alpha_{xy}\propto\varepsilon^{-1}$. Given that $\xi^{2}\propto\varepsilon^{-1}$, this is, indeed, what is expected by Equation 4.
@@ IV.1 Dimensionality of the Gaussian fluctuations
In both equations, the Nernst conductivity is proportional to the quantum of thermoelectric conductance, $k_{b}e/\hbar$, and the only sample dependent parameter is the coherence length. In both equations, the temperature dependence is set by the ratio of $\xi$ and $\ell_{B}$. However, the difference in exponents (reflecting the difference in units (A/K in 2D, and A/K.m in 3D), leads to a qualitatively different temperature dependence: $\alpha_{xy}^{2D}\propto\varepsilon^{-1}$ and $\alpha_{xy}^{3D}\propto\varepsilon^{-1/2}$.
@@
Faced with this apparent paradox, we note first that our situation is neither two-dimensional (the thickness exceeds the coherence length), nor truly three-dimensional (the thickness is not infinite and there is therefore confinement along the c-axis). Now, as seen in equation 3, $\alpha_{xy}$ is the product of the Nernst signal, (which does not depend on dimensionality), and conductivity (which does). The theoretical 3D expression (Equation 7) is derived assuming a stack of 2D layers each $2\xi$ thick. It is the temperature dependence of this thickness which changes the exponent of the power law from -1 (in Equation 4) to -1/2 (in Equation 7). We need to reconcile the relevance of three-dimensionality of $\alpha_{xy}$, which requires the introduction of a relevant length scale along the c-axis, with the existence of a bound on this length scale. Now, since $\xi(T)$ diverges at the critical temperature, the easiest way out is to replace $2\xi(T)$ with $2\xi_{0}$, which has no temperature dependence. In this case, Equation 4 becomes :
@@ V Acknowledgments
VI Data Availability
The supporting data for this article are openly available from the PSL research data repository [58].